OpenStudy (anonymous):

Multiply (sqrt3)(-5sqrt3 +8) use radicals as needed.

OpenStudy (anonymous):

\[\sqrt{3}(-5\sqrt{3}+8)\] \[-15+8\sqrt{3}\]